OpenStudy (anonymous):

60/150 simplest form

OpenStudy (unklerhaukus):

can you factorise 60 ?

OpenStudy (anonymous):

yes?

TheSmartOne (thesmartone):

60:6x10 150:15x10 So now we get 6 and 15 which would make 6/15 Next 6:3x2 15:3x5 So we get 2/5

OpenStudy (anonymous):

60-2*2*3*5

TheSmartOne (thesmartone):

The first time we divided by 10 because 10 was common and the second time 3 is common...

OpenStudy (unklerhaukus):

@qdbug very good , so \[\frac{60}{150}=\frac{2\times2\times3\times5}{150}\] now can you factorise 150

TheSmartOne (thesmartone):

And since 2 and 5 are prime numbers we can not simplify it any furhter...

OpenStudy (anonymous):

TY for the help both of you!

TheSmartOne (thesmartone):

\[\frac{60}{150}=\frac{2\times2\times3\times5}{150}=\frac{ 2 \times 2 \times 3 \times 5 }{ 2 \times 3 \times 5 \times 5 } =\frac{ 2 }{5 }\]